OpenStudy (anonymous):

what is the area of the segment in exact form if the circle has an arc of 120 de

OpenStudy (jdoe0001):

\(\bf \textit{Area of a Circle's Segment}=\cfrac{r^2}{2}\left(\cfrac{\theta\pi}{180}-sin(\theta)\right)\)